OpenStudy (anonymous):

Verify the trig equations match by substituing identities. (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos^(2) x

OpenStudy (anonymous):

\[\sin x(\tan x \cos x-\cot x\cos x)=1-2\cos^2x\\ \sin x\left(\frac{\sin x}{\cos x} \cos x-\frac{\cos x}{\sin x}\cos x\right)=\cdots\] Think you can take it from here?